# Hermite trace polynomials and chaos decompositions for the Hermitian Brownian motion

### v2

@@ 5. The operator algebra
We have the (right) GNS representation of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ on $L^{2}(\mathcal{TP}(\mathcal{H}_{\mathbb{R}}),\varphi)\simeq\mathcal{F}_{q}(\mathcal{H})$, for which the state vector ${(0)}$ is cyclic. For the corresponding representation of $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$, it is cyclic and separating. It follows from Proposition 5.15 that this representation extends to $\overline{\mathcal{TP}}_{q}(\mathcal{H}_{\mathbb{R}})$. Similarly, we have the left representation, which commutes with the right one, and for which $(0)$ is also cyclic.
@@
Let $\mathcal{H}^{\prime}_{\mathbb{R}}\subset\mathcal{H}_{\mathbb{R}}$ be a closed subspace, and $P_{\mathcal{H}^{\prime}}:\mathcal{H}\rightarrow\mathcal{H}^{\prime}$ the orthogonal projection fixing $\mathcal{H}^{\prime}_{\mathbb{R}}$.
@@
The following could have been taken as the definition of $\mathnormal{W}\left(\alpha\otimes F\right)$, with the more explicit expression in terms of contractions obtained a corollary; compare with Section 2.4. Since our definition works on $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$, on which $\varphi$ is not faithful and may not be positive, it is a little more general.
@@
For $q\in\mathcal{Z}$, let $\overline{\mathcal{TP}}_{n}(\mathcal{H}_{\mathbb{R}})$ be the image of $\bigoplus_{k=0}^{n}\mathbb{C}[S_{0}(k)]\otimes\mathcal{H}_{\mathbb{R}}^{\otimes k}$ in $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$. Then the map $\mathnormal{T}\left(\alpha\otimes F\right)\mapsto\mathnormal{W}\left(\alpha\otimes F\right)$ is the orthogonal projection
$\overline{\mathcal{TP}}_{n}(\mathcal{H}_{\mathbb{R}})\rightarrow\overline{\mathcal{TP}}_{n}(\mathcal{H}_{\mathbb{R}})\ominus\overline{\mathcal{TP}}_{n-1}(\mathcal{H}_{\mathbb{R}}).$
@@
It suffices to note that for $\alpha\in S_{0}(k)$, $\mathnormal{W}\left(\alpha\otimes F\right)-\mathnormal{T}\left(\alpha\otimes F\right)\in\overline{\mathcal{TP}}_{k-1}(\mathcal{H}_{\mathbb{R}})$ by Proposition 5.6, and $\mathnormal{W}\left(\alpha\otimes F\right)\perp\overline{\mathcal{TP}}_{k-1}(\mathcal{H}_{\mathbb{R}})$ by definition of the $q$-inner product. ∎
We finish the section with the observation that all the constructions in the article are isomorphic for $q$ and $-q$. In particular, for $q\in\mathcal{Z}$, it suffices to consider $q=\frac{1}{N}$ or $q=0$. We will use the notation $C_{\pi}^{(q)}(\alpha)$, $\mathrm{T}^{(q)}[\alpha\otimes F]$, $\varphi_{q}$, and the multiplication $\cdot_{q}$ to indicate the dependence on $q$.
@@
  • $\mathrm{T}^{(-q)}(R(\alpha\otimes F))=R(\mathrm{T}^{(q)}(\alpha\otimes F))$
@@
    $\varphi_{-q}[R(\mathrm{T}^{(q)}(\alpha\otimes F))]=(-1)^{n}\varphi_{q}[\mathrm{T}^{(q)}(\alpha\otimes F)].$
@@
$\begin{split}\mathrm{T}^{(-q)}(\alpha)&=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}\mathnormal{W}\left(C_{\pi}^{(-q)}(\alpha)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\operatorname{\mathrm{cyc}}_{0}((\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})-\operatorname{\mathrm{cyc}}_{0}(\pi\alpha)+\ell}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{-2\ell-\left|(\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})\right|+\left|\pi\alpha\right|+\ell}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=(-1)^{\left|\alpha\right|}\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{-\ell+\left|\pi\right|}(-1)^{\left|(\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})\right|}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=(-1)^{\left|\alpha\right|}\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|(\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})\right|}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=(-1)^{\left|\alpha\right|}\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}\mathnormal{W}\left(R(C_{\pi}^{(q)}(\alpha))\right)\\ &=(-1)^{\left|\alpha\right|}R(\mathrm{T}^{(q)}(\alpha))\end{split}$
Since
$\begin{split}R(\mathrm{T}^{(q)}(\alpha))\cdot_{-q}R(\mathrm{T}^{(q)}(\beta))&=(-1)^{\left|\alpha\right|+\left|\beta\right|}\mathrm{T}^{(-q)}(\alpha)\cdot_{-q}\mathrm{T}^{(-q)}(\beta)\\ &=(-1)^{\left|\alpha\cup\beta\right|}\mathrm{T}^{(-q)}(\alpha\cup\beta)=R(\mathrm{T}^{(q)}(\alpha)\mathrm{T}^{(q)}(\alpha)),\end{split}$
$R$ is a homomorphism. It clearly commutes with the adjoint operation. Since the action of the state $\varphi$ on $\mathnormal{W}\left(\alpha\otimes F\right)$ does not depend on $q$, and $\left\langle A,B\right\rangle_{q}=\varphi_{q}[B^{\ast}A]$, the isometry property follows from the homomorphism property. Finally, for $\alpha\in S_{0}(2n)$,
$\begin{split}\varphi_{q}\left[(-1)^{\left|\alpha\right|}\mathrm{T}^{(q)}(\alpha\otimes F)\right]&=\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}(-1)^{\left|\alpha\right|}q^{\left|\pi\alpha\right|}C_{\pi}(F)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}(-1)^{\left|\pi\right|}(-q)^{{\pi\alpha}}C_{\pi}(F)\\ &=(-1)^{n}\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}(-q)^{{\pi\alpha}}C_{\pi}(F)\\ &=(-1)^{n}\varphi_{-q}\left[\mathrm{T}^{(-q)}(\alpha\otimes F)\right].\qed\end{split}$
@@ 5.1.1. Gaussian subalgebra
Denote
@@
The distribution of $\mathnormal{W}\left({(0)(1)}\otimes h\right)$ is Gaussian with mean 00 and variance $\left\|h\right\|$.
@@
Thus the distribution of $\mathnormal{W}\left({(0)(1)}\otimes h\right)$ is Gaussian. Finally, since $\mathnormal{T}\left({(0)(1)}\otimes h\right)$ is symmetric, $a^{+}_{(0)(1)}(h)$ maps the $n$’th graded component into the $(n+1)$’st, and $a^{-}_{(0)(1)}(h)$ maps it into the $(n-1)$’st component, these operators have to be each other’s adjoints. ∎
@@
The subspace generated by this algebra’s action on ${(0)}$ is
@@
So the non-degenerate quotient of the space is isomorphic to the symmetric Fock space $\mathcal{F}_{s}(\mathcal{H}_{\mathbb{R}})$, and $\mathcal{TP}_{1}(\mathcal{H}_{\mathbb{R}})=\mathcal{G}(\mathcal{H}_{\mathbb{R}})$.
@@ 5.1.2. Pure trace polynomial subalgebra
$\mathrm{Span}\left(\mathnormal{W}\left(\alpha\otimes_{s}F\right):\alpha(0)=0\right)$
is the center of $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$.
@@
Note that
@@
If $\alpha\in S(n)$, $\beta\in S_{0}(k)$, then $\beta$ and $\sigma_{n,k}\alpha\sigma_{n,k}^{-1}$ commute, and the expression above is $\beta\cup\alpha$. Since also $U_{\sigma_{n,k}}(F\otimes G)=G\otimes F$,
@@
For the converse, suppose that for some for $A\in\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$, $A\ \mathnormal{W}\left((01)\otimes h\right)=\mathnormal{W}\left((01)\otimes h\right)A$ for all $h\in\mathcal{H}$. $A$ has the form
$A=\sum_{k=0}^{n}\sum_{\alpha\in S_{0}(k)}\mathnormal{W}\left(\alpha\otimes_{s}F_{\alpha}\right).$
It suffices to show that for each $\alpha\in S_{0}(n)$, $\alpha(0)\neq 0$ implies that $F_{\alpha}=0$ (since such a term can be subtracted from $A$ with the difference still in the center). Comparing only the terms in the $(n+1)$’th component, it follows that
$\sum_{\alpha\in S_{0}(n)}\mathnormal{W}\left((\alpha\cup(01))\otimes_{s}(F_{\alpha}\otimes h)\right)=\sum_{\beta\in S_{0}(n)}\mathnormal{W}\left(((01)\cup\beta)\otimes_{s}(h\otimes F_{\beta})\right).$
@@
$\sum_{\alpha\in S_{0}(n)}\mathnormal{W}\left(((0\ n+1)\alpha)\otimes_{s}(F_{\alpha}\otimes h)\right)=\sum_{\beta\in S_{0}(n)}\mathnormal{W}\left((\beta(0\ n+1))\otimes_{s}(F_{\beta}\otimes h)\right).$
Suppose $F_{\alpha}\neq 0$, and for some $\sigma\in S(n+1)$,
$(0\ n+1)\alpha=\sigma\beta(0\ n+1)\sigma^{-1}\text{ and }F_{\alpha}\otimes h=U_{\sigma}(F_{\beta}\otimes h).$
Then from the second relation, in fact $\sigma\in S(n)$. Applying the first relation to 00, it follows that $\alpha(0)=0$. ∎

### v3

@@ 5. The operator algebra
Let $q\in\mathcal{Z}$,. By Theorem 5.15(a,b,c), the GNS Hilbert space $L^{2}(\mathcal{TP}(\mathcal{H}_{\mathbb{R}}),\varphi)$ is isomorphic to $\mathcal{F}_{q}(\mathcal{H})$, and have the (right) star-representation of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ on this Hilbert space, for which the state vector (0){(0)} is cyclic, and $\mathnormal{W}\left(\alpha\otimes_{s}F\right)(0)=\alpha\otimes_{s}F\in\mathcal{F}_{q}(\mathcal{H})$. For the corresponding representation of $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$, it is cyclic and separating. It follows from Proposition 5.16 that this representation extends to $\overline{\mathcal{TP}}_{q}(\mathcal{H}_{\mathbb{R}})$. Similarly, we have the left representation, which commutes with the right one, and for which $(0)$ is also cyclic.
@@
Let $\mathcal{H}^{\prime}_{\mathbb{R}}\subset\mathcal{H}$$_{\mathbb{R}}$ be a closed subspace, let $\mathcal{H}^{\prime}=\mathbb{C}\mathcal{H}_{\mathbb{R}}$$^{\prime}$, and let $P_{\mathcal{H}^{\prime}}$ be the orthogonal projection onto $\mathcal{H}^{\prime}$. Fix $q\in\mathcal{Z}$.
@@
The following could have been taken as the definition of $\mathnormal{W}\left(\alpha\otimes F\right)$, with the more explicit expression in terms of contractions obtained a corollary; compare with Section 2.4. Since our algebraic definition works on $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$, on which $\varphi$ is not faithful and may not be positive, it is a little more general.
@@
For $q\in\mathcal{Z}$, let $\overline{\mathcal{TP}}_{n}(\mathcal{H}_{\mathbb{R}})$ be the image of $\bigoplus_{k=0}^{n}\mathbb{C}[S_{0}(k)]\otimes\mathcal{H}_{\mathbb{R}}^{\otimes k}$ in $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$. Denote by $P_{n}$ the orthogonal projection
$P_{n}:\overline{\mathcal{TP}}_{n}(\mathcal{H}_{\mathbb{R}})\rightarrow\overline{\mathcal{TP}}_{n}(\mathcal{H}_{\mathbb{R}})\ominus\overline{\mathcal{TP}}_{n-1}(\mathcal{H}_{\mathbb{R}}).$
Then for $\alpha\otimes F\in\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}_{\mathbb{R}}^{\otimes n}$,
$\mathnormal{W}\left(\alpha\otimes F\right)=P_{n}\mathnormal{T}\left(\alpha\otimes F\right)$
@@
It suffices to note that for $\alpha\in S_{0}(n)$, $\mathnormal{W}\left(\alpha\otimes F\right)-\mathnormal{T}\left(\alpha\otimes F\right)\in\overline{\mathcal{TP}}_{n-1}(\mathcal{H}_{\mathbb{R}})$ by Proposition 5.7, and $\mathnormal{W}\left(\alpha\otimes F\right)\perp\overline{\mathcal{TP}}_{n-1}(\mathcal{H}_{\mathbb{R}})$ by definition of the $q$-inner product. ∎
We finish the section with the observation that all the constructions in the article are isomorphic for $q$ and q-q. In particular, for $q\in\mathcal{Z}$, it suffices to consider $q=\frac{1}{N}$ or $q=0$. We will use the notation $C_{\pi}^{(q)}(\alpha)$, $T^{(q)}[\alpha\otimes F]$, $\varphi_{q}$, and the multiplication $\cdot_{q}$ to indicate the dependence on $q$.
@@
  • $T^{(-q)}(R(\alpha\otimes F))=R(T^{(q)}(\alpha\otimes F))$
@@
    $\varphi_{-q}[R(T^{(q)}(\alpha\otimes F))]=(-1)^{n}\varphi_{q}[T^{(q)}(\alpha\otimes F)].$
@@
$\begin{split}T^{(-q)}(\alpha)&=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}\mathnormal{W}\left(C_{\pi}^{(-q)}(\alpha)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\operatorname{\mathrm{cyc}}_{0}((\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})-\operatorname{\mathrm{cyc}}_{0}(\pi\alpha)+\left|\pi\right|}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left(n-2\left|\pi\right|-\left|(\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})\right|\right)-(n-\left|\pi\alpha\right|)+\left|\pi\right|}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|\pi\right|+\left|\pi\alpha\right|}(-1)^{\left|(\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})\right|}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=(-1)^{\left|\alpha\right|}\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|(\pi\alpha)|_{\mathrm{supp}\left(\pi\right)^{c}})\right|}\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\\ &=(-1)^{\left|\alpha\right|}\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}R\left(\mathnormal{W}\left(C_{\pi}^{(q)}(\alpha)\right)\right)\\ &=(-1)^{\left|\alpha\right|}R(T^{(q)}(\alpha))\end{split}.$
Here we used the fact that for any $\alpha,\pi$, $\left|\alpha\right|+\left|\pi\alpha\right|+\left|\pi\right|$ is even. Since
$\begin{split}R(T^{(q)}(\alpha))\cdot_{-q}R(T^{(q)}(\beta))&=(-1)^{\left|\alpha\right|+\left|\beta\right|}T^{(-q)}(\alpha)\cdot_{-q}T^{(-q)}(\beta)\\ &=(-1)^{\left|\alpha\cup\beta\right|}T^{(-q)}(\alpha\cup\beta)=R(T^{(q)}(\alpha)T^{(q)}(\alpha)),\end{split}$
$R$ is a homomorphism. It clearly commutes with the adjoint operation. Since the action of the linear functional $\varphi$ on $\mathnormal{W}\left(\alpha\otimes F\right)$ does not depend on $q$, and $\left\langle A,B\right\rangle_{q}=\varphi_{q}[B^{\ast}A]$, the isometry property follows from the homomorphism property. Finally, for $\alpha\in S_{0}(2n)$,
$\begin{split}\varphi_{q}\left[(-1)^{\left|\alpha\right|}T^{(q)}(\alpha\otimes F)\right]&=\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}(-1)^{\left|\alpha\right|}q^{\left|\pi\alpha\right|}C_{\pi}(F)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}(-1)^{\left|\pi\right|}(-q)^{{\pi\alpha}}C_{\pi}(F)\\ &=(-1)^{n}\sum_{\pi\in\operatorname{\mathcal{P}}_{2}(2n)}(-q)^{{\pi\alpha}}C_{\pi}(F)\\ &=(-1)^{n}\varphi_{-q}\left[{}^{(-q)}(\alpha\otimes F)\right].\qed\end{split}$
@@ 5.1.1. Gaussian subalgebra
Definition 5.23.
The Gaussian subalgebra of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ is
@@
The distribution of $\mathnormal{W}\left({(0)(1)}\otimes h\right)$ is Gaussian with mean 0 and variance $\left\|h\right\|$.
@@
Thus the distribution of $\mathnormal{W}\left({(0)(1)}\otimes h\right)$ is Gaussian. Finally, since $\mathnormal{T}\left({(0)(1)}\otimes h\right)$ is symmetric, $a^{+}_{(0)(1)}(h)$ maps the $n$’th graded component into the $(n+1)$’st, and $a^{-}_{(0)(1)}(h)$ maps it into the (n1)(n-1)’st component, these operators have to be each other’s adjoints. ∎
@@
The subspace generated by this algebra’s action on (0){(0)} is
@@
So the non-degenerate quotient of the space is isomorphic to the symmetric Fock space $\mathcal{F}_{s}(\mathcal{H})$, and $\mathcal{TP}_{1}(\mathcal{H}_{\mathbb{R}})=\mathcal{G}(\mathcal{H}_{\mathbb{R}})$.
@@ 5.1.2. Pure trace polynomial subalgebra
Definition 5.27.
The pure trace polynomial subalgebra of $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ is
$\mathrm{Span}\left(\mathnormal{W}\left(\alpha\otimes_{s}F\right):\alpha(0)=0\right).$
Remark 5.28.
Recall that
$\overline{\mathcal{TP}}(\mathbb{C})=\bigoplus_{n=0}^{\infty}Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])$
is commutative.
@@
Let $\dim\mathcal{H}_{\mathbb{R}}\geq 2$. Then the pure trace polynomial subalgebra is the center of $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$.
@@
Recall from Notation 3.2 that
@@
If $\alpha\in S(n)$ and $\beta\in S_{0}(k)$, then $\beta$ and $\sigma_{n,k}\alpha\sigma_{n,k}^{-1}$ commute, and the expression above is $\beta\cup\alpha$. Since also $U_{\sigma_{n,k}}(F\otimes G)=G\otimes F$,
@@
For the converse, denote $\left\{\xi_{i}:i\in\Xi\right\}$ an orthonormal basis for $\mathcal{H}_{\mathbb{R}}$. Suppose that for some $A\in\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$, $A\ \mathnormal{W}\left((01)\otimes\xi_{i}\right)=\mathnormal{W}\left((01)\otimes\xi_{i}\right)A$ for all $i\in\Xi$. $A$ has the form
$A=\sum_{n=0}^{k}\sum_{\alpha\in S_{0}(n)}\mathnormal{W}\left(\alpha\otimes F_{\alpha}\right),$
where for each $n$ and $F_{\alpha}\in\mathcal{H}_{\mathbb{R}}^{\otimes n}$,
(14) $\frac{1}{n!}\sum_{\sigma\in S(n)}U_{\sigma}F_{\sigma^{-1}\alpha\sigma}=F_{\alpha}.$
Comparing only the terms in the $(n+1)$’th component, it follows that
$\sum_{\alpha\in S_{0}(n)}\mathnormal{W}\left((\alpha\cup(01))\otimes_{s}(F_{\alpha}\otimes\xi_{i})\right)=\sum_{\beta\in S_{0}(n)}\mathnormal{W}\left(((01)\cup\beta)\otimes_{s}(\xi_{i}\otimes F_{\beta})\right).$
@@
$\sum_{\alpha\in S_{0}(n)}\mathnormal{W}\left(((0\ n+1)\alpha)\otimes_{s}(F_{\alpha}\otimes\xi_{i})\right)=\sum_{\beta\in S_{0}(n)}\mathnormal{W}\left((\beta(0\ n+1))\otimes_{s}(F_{\beta}\otimes\xi_{i})\right).$
That is,
(15) $\sum_{\sigma\in S(n+1)}\sum_{\alpha\in S_{0}(n)}\mathnormal{W}\left(\sigma((0\ n+1)\alpha)\sigma^{-1}\otimes U_{\sigma}(F_{\alpha}\otimes\xi_{i})\right)\\ =\sum_{\tau\in S(n+1)}\sum_{\beta\in S_{0}(n)}\mathnormal{W}\left(\tau(\beta(0\ n+1))\tau^{-1}\otimes U_{\tau}(F_{\beta}\otimes\xi_{i})\right).$
Fix $\tilde{\alpha}\in S_{0}(n)$ with $\tilde{\alpha}(0)\neq 0$, and denote $\gamma=(0\ n+1)\tilde{\alpha}$. Note that if $\gamma=\sigma(0\ n+1)\alpha\sigma^{-1}$ with $\alpha\in S_{0}(n)$ and $\sigma\in S(n+1)$, then in fact $\sigma\in S(n)$. This follows since
$\sigma(n+1)=(\tilde{\alpha}^{-1}(0\ n+1)\sigma(0\ n+1)\alpha)(n+1)=n+1.$
Next, from (15),
$\sum_{\begin{subarray}{c}\sigma\in S(n+1),\alpha\in S_{0}(n)\\ \sigma(0\ n+1)\alpha\sigma^{-1}=\gamma\end{subarray}}U_{\sigma}(F_{\alpha}\otimes\xi_{i})=\sum_{\begin{subarray}{c}\tau\in S(n+1),\beta\in S_{0}(n)\\ \tau\beta(0\ n+1)\tau^{-1}=\gamma\end{subarray}}U_{\tau}(F_{\beta}\otimes\xi_{i}).$
The left-hand simplifies to
$\sum_{\begin{subarray}{c}\sigma\in S(n),\alpha\in S_{0}(n)\\ \sigma\alpha\sigma^{-1}=\tilde{\alpha}\end{subarray}}U_{\sigma}(F_{\alpha}\otimes\xi_{i})=\sum_{\sigma\in S(n)}U_{\sigma}(F_{\sigma^{-1}\tilde{\alpha}\sigma}\otimes\xi_{i})=n!F_{\tilde{\alpha}}\otimes\xi_{i},$
where in the last step we used (14). Note also that for $\tau,\beta$ as in the sum above,
(16) $\gamma(0)=\tilde{\alpha}(0)=\tau(n+1),$
where we recall that $\tilde{\alpha}(0)\neq 0,n+1$. Therefore
$n!F_{\tilde{\alpha}}\otimes\xi_{i}=\sum_{\begin{subarray}{c}\tau\in S(n+1),\beta\in S_{0}(n)\\ \tau\beta(0\ n+1)\tau^{-1}=\gamma\end{subarray}}U_{\tau}(F_{\beta}\otimes\xi_{i})\in U_{(\tilde{\alpha}(0)\ n+1)}(\mathcal{H}_{\mathbb{R}}^{\otimes(n-1)}\otimes\mathrm{Span}\left(\xi_{i}\right)).$
Since this is true for each $i\in\Xi$, which has at least 2 elements, it follows that $F_{\tilde{\alpha}}=0$. ∎